# Hypergraph Representations of scRNA-seq Data for Improved Clustering with Random Walks

### v2

@@ III.3 Sparsity-Induced Correlation Inflation in scRNA-seq Co-expression Networks
The zero inflation problem in single-cell RNA sequencing (scRNA-seq) data analysis refers to the excessive number of zero counts (or "dropouts") observed in the data. These zeros could either be a result of technical limitation in detecting low abundance transcripts (false negatives), or true absence biological zeros (true negatives) [37, 39, 40]. In scRNA-seq data, when individually profiled cell expression are merged to create a unified dataset, all genes are included in the expression profile of every cell. However, since cells typically express different sets of genes, the inclusion of all genes introduces significant sparsity into the merged dataset. Consequentially, when the scRNA-seq data are analyzed through co-expression networks, these induced zeros can inflate correlation coefficients between cells or genes. This occurs because during co-expression network construction, edge weights represent correlation between expression profiles and are often computed by Pearson’s correlation coefficient or Spearman’s rank correlation coefficient. In these commonly adopted correlation measures, zeros are treated as expression magnitudes in the same way as non-zeros. As a result, if two expression profiles share a large number of common zeros, their correlation will be inflated since the common zeros will be considered as evidence of their expression homogeneity, which can lead to an overestimation of the similarity between expression profiles. This issue is particularly pronounced at the single-cell level compared to bulk tissue RNAseq data, as the proportion of genes expressed at the single-cell level is much smaller than at the tissue level, resulting in a higher degree of induced data sparsity.
In the experiment shown in Fig. 3, we investigate the impact of induced sparsity in scRNA-seq data on correlation coefficients and its implications for interpreting gene and cell connections in co-expression networks. We demonstrate that induced zeros lead to inflated correlation coefficients between cells or genes. To simulate this effect, we generated gene expression data with a fixed base size and appended varying proportions of zeros to mimic different degrees of sparsity induced when merging cell expression profiles. We show that as the proportion of zeros appended to the data increases, the correlation coefficient also increases, leading to an overestimation of the similarity between expression profiles. Our results show that as the level of induced sparsity increases, the correlation coefficient also increases, leading to an overestimation of similarity between expression profiles. These findings highlight the importance of considering zero inflation in scRNA-seq data analysis and the need for accurate representation of the underlying data structures.
@@ III.4.1 Evaluation Metrics
III.4.2 Clustering Performance vs. Bipartite Modularity
We further evaluated the performance of these clustering methods under varying data modularity conditions. The modularity variation was controlled by two key parameters: the average number of co-expressed genes per cell type and the number of embedded cell modules. Additionally, we implemented Barber’s bipartite modularity [42], as shown by the barplots in Fig.4, to help quantify modularity of the underlying data for clustering and analyze the association between modularity and clustering performance. As shown in Fig.4, clustering performance generally improves with increased data modularity. The modularity bar plots in Fig.4 and heatmap visualization in Fig.S6 demonstrate that the simulated scRNA-seq data displays stronger modularity when the average number of co-expressed genes per cell type is greater or when the number of modules is smaller. Our proposed methods—DIPHW, CoMem, and CoMem-DIPHW—consistently outperformed all other clustering methods that use unipartite network projections, including Node2Vec, and community detection-based methods such as Louvain and Greedy Modularity. Notably, CoMem-DIPHW consistently demonstrated superior performance across almost all parameter settings, with its advantage being particularly pronounced under weak data modularity conditions.
@@
III.4.3 Clustering Performance Comparison under strong and weak data modularity parameter regimes
To evaluate clustering performance of the proposed methods across modularity conditions, we considered two parameters that determine the modularity of the simulated scRNA-seq data: (i) the average number of co-expressed genes per module and (ii) the number of embedded modules. The number of modules here represents the number of cell types in the simulated scRNA-seq data. Weak modularity corresponds to a small number of co-expressed genes per module or a large number of embedded modules, making data structures more challenging to detect. Conversely, strong modularity corresponds to a large number of co-expressed genes per module or a small number of embedded modules. For clarity of presentation, we present the most contrasting cases from our analysis: (i) an average of 10 and 100 co-expressed genes per module in Fig. 5 and (ii) 10 and 100 embedded modules in Fig. 6. In Fig. 5 and Fig. 6, we compare the performance of our proposed hypergraph-based clustering methods (highlighted in red and yellow) against five embedding-based methods and four community detection methods under both strong and weak modularity conditions. All embedding-based methods are coupled with K-means to produce clustering results. In weak modularity regimes (where the average number of co-expressed genes = 10 or when number of modules = 100), CoMem-DIPHW and DIPHW exhibited a significant performance advantage—for example, when the average number of co-expressed genes = 10, CoMem-DIPHW had an ARI of 0.29 compared to 0.05 for the best-performing baseline method (PCA), representing a performance improvement of approximately 450%. Under strong modularity conditions (when the average number of genes per module = 100 or when number of modules = 10), several methods—including PCA, CoMem-DIPHW, CoMem, Node2Vec, and DIPHW—performed well, achieving average ARI scores above 0.9.
@@ III.4.3 Clustering Performance Comparison under strong and weak data modularity parameter regimes
III.4.4 Evaluating the Impact of Correlation Methods on Network-Based Clustering Performance
In Table 3, we evaluated how co-expression networks generated by different correlation computation methods lead to varying performance across various network-based clustering methods. The correlation methods considered were Pearson’s, Spearman’s, Kendall’s Tau, and cosine similarity. Our results show that cosine similarity often led to the worst clustering results among the correlation computation methods tested.
Method ARI NMI
Pearson Cosine Spearman Kendall Pearson Cosine Spearman Kendall
GreedyModularity 0.205 0.135 0.194 0.197 0.534 0.429 0.496 0.502
Louvain 0.493 0.322 0.445 0.473 0.784 0.666 0.768 0.781
Infomap 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000
Leiden -0.001 -0.001 -$0$0.000 0.062 0.040 0.082 0.073
Multilevel 0.032 0.034 0.034 0.035 0.305 0.289 0.325 0.319
Eigenvector 0.002 0.000 0.003 0.005 0.097 0.060 0.082 0.091
Table 3: Impact of Correlation Methods on Network-Based Clustering Performance. The table evaluates the impact of correlation computation methods—Pearson’s, Spearman’s, Kendall’s Tau, and cosine similarity—on the performance of various network-based clustering methods. The correlation computation method resulting in the worst clustering performance is marked in bold. Cosine similarity, which ignores shared zeros in cell expression profiles, often resulted in the worst clustering performance. These results show that, although using correlation computation methods that treat zeros the same as non-zero signals could inflate correlations, ignoring these zeros, as in cosine similarity, can lead to worse performance.
@@ III.4.4 Evaluating the Impact of Correlation Methods on Network-Based Clustering Performance
A major challenge in scRNA-seq analysis is the lack of reliable ground-truth for cell type annotation, which complicates the evaluation of clustering algorithms. In most scRNA-seq datasets, cell type labels are typically assigned using a combination of clustering results, differential expression (DE) analysis, and canonical marker genes. First, cells are grouped into clusters based on transcriptomic similarity using clustering algorithms (e.g., Leiden, Louvain, or k-means). Then, DE analysis is performed to identify genes that are differentially expressed across clusters. Finally, these clusters are annotated based on known canonical marker genes associated with specific cell types. Because this annotation process relies on clustering results, using these labels to evaluate clustering performance introduces circular logic and bias toward the method originally used for annotation. To address this issue, we utilized the mixture control benchmarking dataset [61], where ground truth is established based on cell line identity.
@@ III.6 Human Pancreas Results
We report the top 10 differentially expressed genes (DEGs) from each human pancreas cell cluster identified by CoMem-DIPHW and PCA in Fig. 7a and Fig. 8a respectively. To evaluate the clustering performance of CoMem-DIPHW applied to human pancreas scRNA-seq data, we assessed the across-cluster differentiation in the average expression of the cluster-specific DEGs, as shown in Fig. 7b. Strong clustering performance is indicated by high expression of cluster-specific DEGs within their respective clusters and low expression in other clusters. To compare clustering performance between PCA and CoMem-DIPHW, we examined the expression patterns of cluster-specific DEGs identified by CoMem-DIPHW (Fig. 7b) and PCA (Fig. 8b). In these figures, each subplot corresponds to one cluster’s DEGs, with violin plots showing their average expression levels across all clusters. The DEGs identified by CoMem-DIPHW exhibit more cluster-specific expression patterns compared to DEGs found by PCA. For example, comparing the PCA clustering results in Fig. 8b, the average expression of the DEGs from PCA cluster 1 is higher in clusters 2, 5, and 6. Similarly, the DEGs from cluster 7 show higher average expression in cluster 6. In general, the average expression of the DEGs from the PCA clusters is less cluster-specific than the DEGs from the CoMem-DIPHW clusters.
@@
Moreover, we used the canonical gene markers reported in the original study of the human pancreas scRNA-seq data [58] to tentatively label the cell clusters found by CoMem-DIPHW. The markers include GCG (alpha cells), INS (beta cells), SST (delta cells), PPY (PP cells), PRSS1 (acinar cells), KRT19 (ductal cells), COL1A1 (mesenchymal cells), and ESAM (endothelial cells). Clusters with high expression of corresponding gene markers are assigned to their respective cell types. As shown in Fig. 7d, each cell-type specific marker was distinctly expressed in a cohesive and well-defined cluster in the embedding generated by CoMem-DIPHW. The results show that CoMem-DIPHW embeddings effectively distinguish different cell types. In contrast, as shown in Fig. 8d, PCA embeddings lacked clear separation between different cell types.
As shown in Fig. 7d, Cluster 0 was identified as acinar cells by the high expression of PRSS1. In Table 7a, the top 10 DEGs for cluster 0 include 5 genes involved in digestive enzyme production, a primary function of pancreatic acinar cells: CPA1, CPA2 [68], PLA2G1B [69], CTRB1 [70], and PRSS1 [71]. Clusters 2 and 7 were identified as alpha cells by the high expression of GCG. Alpha cells are responsible for glucagon secretion and blood glucose regulation [72, 73]. Markers for alpha cells among the DEGs in cluster 2 found in the PanglaoDB database include GLS, CRYBA2, FAP, LOXL4, MAFB, and RGS4. In cluster 7, alpha cell marker DEGs include LOXL4, PAX6, RGS4, and CRYBA2. Cluster 3 was classified as ductal cells by the high expression of KRT19. Six out of the 10 DEGs identified by CoMem-DIPHW are markers for ductal cells in the pancreas, including SLC4A4, ANXA4, KRT19, CFTR, CLDN1, and ONECUT2. Cluster 5 was classified as delta cells by the high expression of SST. Established markers strongly associated with delta cells in DEGs identified by CoMem-DIPHW include GAD2, PCSK1, LEPR, and CASR. Cluster 6 was classified as beta cells by the high expression of INS. This is also strongly supported by the DEGs MAFA, ERO1LB, PDX1, IAPP, SYT13, HADH and ADCYAP1. Both COL1A1 and ESAM are highly expressed in cluster 4, suggesting mesenchymal or endothelial cells. These align well with 9 out of the top 10 DEGs found by CoMem-DIPHW, including COL4A2, COL4A1, SPARC, EMP1, IGFBP4, HTRA1, PXDN, LAMA4 and NID1.
@@
To further validate the tentative cell type annotations and clustering results, we cross-referenced the cluster-specific DEGs (see Fig.7a) with markers from the PanglaoDB cell type marker database (version 27 Mar 2020)[74]. The assignment of cell types was determined based on the overlap between the DEGs of each cluster and the cell type-specific markers in the PanglaoDB marker database, with a match score computed based on the proportion of matched markers. Cell types with a score above a threshold of 0.05 were considered potential matches. Table 7c presents the top matches for each cluster. These results align well with our previous analysis based on marker gene distribution (Figure 7d), providing additional support for our cell type assignments. The only discrepancy between the results based on PanglaoDB and the marker gene distribution experiments lies in cluster 4. Matching the DEGs with PanglaoDB cell type markers suggests Pancreatic stellate cells (PSCs) or Granulosa cells, while the COL1A1 and ESAM expression plots indicate mesenchymal or endothelial cells. However, this apparent discrepancy can be explained by the fact that PSCs can differentiate into cells with endothelial-like properties [75], and granulosa cells can exhibit mesenchymal-like behaviors [76]. Both cell types share characteristics with mesenchymal and endothelial cells, explaining the overlap in gene expression patterns observed in cluster 4. Hence, the cell clusters determined by CoMem-DIPHW and annotated by the canonical markers selected by the original study [58] are also supported by the strong overlap between each cluster’s top DEGs and cell type-specific markers from the PanglaoDB database.
@@
On the other hand, the differential analysis based on the PCA clustering results on the human pancreas dataset (Fig.8) shows limited cluster separation. While some of the PCA clusters could be consistently annotated with cell types based on evidence from both PanglaoDB markers and canonical markers, considerable overlap remains among PCA clusters. For example, cluster 8 is identified as beta cells by both the PanglaoDB markers and the canonical marker INS. However, cluster 8 is not well separated from cluster 5, which potentially represents delta cells based on its high expression of SST. Similarly, cluster 7 is identified as ductal cells by both the PanglaoDB markers and the canonical marker KRT19, while the endothelial cells marked by the high expression of ESAM are embedded very close to cluster 7 in the PCA embedding.
@@
Figure 7: CoMem-DIPHW Clustering Performance Assessment on the Humen Pancreas Dataset and Cell Type Annotation using Differentially Expressed Genes (DEGs) and Canonical Markers. (a) Top 10 DEGs identified for each cluster by CoMem-DIPHW. (b) Across-cluster average expression of cluster-specific DEGs. Violin plots show the distribution of average expression levels of these DEGs across all clusters. Strong clustering performance is indicated by high expression of cluster-specific DEGs within their respective clusters and low expression in other clusters. (c) Cell type annotation using the PanglaoDB marker database. Cell types are determined by the overlap between each cluster’s DEGs and cell type-specific markers in the PanglaoDB database, with match scores computed based on the proportion of matched markers. (d) Expression of human pancreatic canonical markers in CoMem-DIPHW embeddings. UMAP visualizations show CoMem-DIPHW cell embeddings, with cluster IDs assigned by K-means clustering. Each subplot displays the expression of a canonical marker gene for a specific human pancreatic cell type. The cell-type specific markers considered include: GCG (alpha cells), INS (beta cells), SST (delta cells), PPY (PP cells), PRSS1 (acinar cells), KRT19 (ductal cells), COL1A1 (mesenchymal cells), and ESAM (endothelial cells). Color intensity represents gene expression levels, with redder hues indicating higher expression. Clusters show high expression of cell-type-specific canonical markers are assigned to the corresponding cell types. Overall, the clustering quality is demonstrated through: high within-cluster expression of DEGs shown in the violin plots in (b), distinct separation of cell types in (d), and consistency between canonical marker-based annotation in (d) and PanglaoDB-based annotation in (c).
@@
Figure 8: PCA Performance Assessment on the Humen Pancreas Dataset and Cluster Cell Type Annotation using DEGs and Canonical Markers. (a) Top 10 DEGs identified for each cluster by PCA. (b) Across-cluster average expression of cluster-specific DEGs. (c) Cell type annotation using the PanglaoDB marker database. (d) Expression of human pancreatic canonical markers in PCA embeddings. The cell-type specific markers considered include: GCG (alpha cells), INS (beta cells), SST (delta cells), PPY (PP cells), PRSS1 (acinar cells), KRT19 (ductal cells), COL1A1 (mesenchymal cells), and ESAM (endothelial cells). UMAP visualizations show PCA cell embeddings, with cluster IDs assigned by K-means clustering. The overall cluster or cell type separation is limited compared to CoMem-DIPHW clustering, as evidenced by the overlapping beta cells, delta cells and PP cells in (d) and the lack of distinct high expression patterns of the DEGs in (b).
@@
Following discussion of the results from human pancreas dataset, we provide a brief overview of the results from the mouse pancreas, human brain, and mouse brain datasets. For these datasets, we have included visualizations in the appendix that show the average expression of the top DEGs across clusters identified by different methods, which shows our proposed algorithm’s ability in effectively distinguishing cell clusters by their expression profiles. The supplements also contain tables listing the top DEGs for clusters identified by CoMem-DIPHW and PCA in each dataset for reference.

### v3

@@ III.3 Sparsity-Induced Correlation Inflation in scRNA-seq Co-expression Networks
The zero inflation problem in single-cell RNA sequencing (scRNA-seq) data analysis refers to the excessive number of zero counts (or "dropouts") observed in the data. These zeros could either be a result of technical limitation in detecting low abundance transcripts (false negatives), or true absence of biological zeros (true negatives) [37, 39, 40]. The process of integrating individual cell expression profiles into a unified scRNA-seq dataset requires constructing a matrix that accounts for every gene detected in any cell, which artificially introduces zero entries wherever genes are not expressed in those cells. Since cells typically express different sets of genes, this inclusion introduces significant sparsity into the merged dataset. In co-expression network construction, edge weights are inflated due to these artificially introduced common zeros since the common zeros will be considered as evidence of their expression homogeneity. This issue is particularly pronounced at the single-cell level compared to bulk tissue RNAseq data, as the proportion of expressed genes at the single-cell level is much smaller, resulting in higher induced data sparsity.
In the experiment shown in Fig. 3, we investigate how induced sparsity affects correlation coefficients and the interpretation of gene and cell connections in co-expression networks. To simulate this effect, we generated gene expression data with a fixed base size and introduced varying proportions of zeros to mimic different sparsity levels. Our results show that as the level of induced sparsity increases, correlation coefficients between expression profiles become artificially inflated, leading to an overestimation of similarity between cell expression profiles. These findings highlight the importance of addressing zero inflation in scRNA-seq data analysis and the need for accurate representation of the underlying biological relationships. While induced sparsity inflates correlations, our supplementary experiments (.3) show that ignoring shared zeros using cosine similarity leads to worse clustering performance, as these inactive genes still hold biological significance.
@@ III.4.1 Evaluation Metrics
III.4.2 Clustering Performance vs. Bipartite Modularity and Method Comparison
We evaluated the performance of these clustering methods under varying data modularity conditions. The modularity variation was controlled by two key parameters: the average number of co-expressed genes per cell type and the number of embedded cell modules. The number of modules here represents the number of cell types in the simulated scRNA-seq data. Weak modularity corresponds to a small number of co-expressed genes per module or a large number of embedded modules, making data structures more challenging to detect. The modularity bar plots in Fig. 4 and heatmap visualization in Fig. S6 demonstrate that the simulated scRNA-seq data displays stronger modularity when the average number of co-expressed genes per cell type is greater or when the number of modules is smaller. We implemented Barber’s bipartite modularity [42], as shown by the barplots in Fig. 4, to help quantify modularity of the underlying data for clustering and analyze the association between modularity and clustering performance. Clustering performance generally improves with increased data modularity. Our proposed methods—DIPHW, CoMem, and CoMem-DIPHW—consistently outperformed all other clustering methods that use unipartite network projections, including Node2Vec, and community detection-based methods such as Louvain and Greedy Modularity. Notably, CoMem-DIPHW consistently demonstrated superior performance across almost all parameter settings, with its advantage being particularly pronounced under weak data modularity conditions. For clarity of presentation, we present the most contrasting cases from our analysis: (i) an average of 10 and 100 co-expressed genes per module in Fig. 5 and (ii) 10 and 100 embedded modules in Fig. 6.
@@
In Fig. 5 and Fig. 6, we compare the performance of our proposed hypergraph-based clustering methods (highlighted in red and yellow) against five embedding-based methods and four community detection methods under both strong and weak modularity conditions. All embedding-based methods are coupled with K-means to produce clustering results. In weak modularity regimes (where the average number of co-expressed genes = 10 or when number of modules = 100), CoMem-DIPHW and DIPHW exhibited a significant performance advantage—for example, when the average number of co-expressed genes = 10, CoMem-DIPHW had an ARI of 0.29 compared to 0.05 for the best-performing baseline method (PCA), representing a performance improvement of approximately 450%. Under strong modularity conditions (when the average number of genes per module = 100 or when number of modules = 10), several methods—including PCA, CoMem-DIPHW, CoMem, Node2Vec, and DIPHW—performed well, achieving average ARI scores above 0.9.
@@
A major challenge in scRNA-seq analysis is the lack of reliable ground-truth for cell type annotation, which complicates the evaluation of clustering algorithms. In most scRNA-seq datasets, cell type labels are typically assigned using a combination of clustering results, differential expression (DE) analysis, and canonical marker genes. First, cells are grouped into clusters based on transcriptomic similarity using clustering algorithms (e.g., Leiden, Louvain, or k-means). Then, DE analysis is performed to identify genes that are differentially expressed across clusters. Finally, these clusters are annotated based on known canonical marker genes associated with specific cell types. Since this annotation process relies on clustering results, using these labels to evaluate clustering performance introduces circular logic and bias toward the method originally used for annotation. To address this issue, we utilized the mixture control benchmarking dataset [61], where ground truth is established based on cell line identity.
@@ III.6 Human Pancreas Results
We report the top 10 differentially expressed genes (DEGs) from each human pancreas cell cluster identified by CoMem-DIPHW and PCA in Fig. 7a and Fig. 8a respectively. To evaluate clustering performance, we examined the expression of these DEGs across clusters (Fig. 7b) and PCA (Fig. 8b). Violin plots show that the DEGs identified by CoMem-DIPHW exhibit more cluster-specific expression patterns than PCA-derived DEGs, which tend to show higher expression in multiple clusters. For example, the average expression of the DEGs from PCA cluster 1 is higher in clusters 2, 5, and 6. Similarly, the DEGs from cluster 7 show higher average expression in cluster 6. In general, the average expression of the DEGs from the PCA clusters is less cluster-specific than the DEGs from the CoMem-DIPHW clusters.
@@
To tentatively label the cell clusters, we used canonical gene markers reported in the original human pancreas scRNA-seq study [58]. We assigned clusters based on the expression of these canonical markers including GCG (alpha cells), INS (beta cells), SST (delta cells), PPY (PP cells), PRSS1 (acinar cells), KRT19 (ductal cells), COL1A1 (mesenchymal cells), and ESAM (endothelial cells). Each marker is distinctly expressed in a well-defined cluster in the embedding generated by CoMem-DIPHW, as shown in Fig. 7d, confirming that CoMem-DIPHW embeddings effectively distinguish different cell types. In contrast, PCA embeddings lacked clear separation between different cell types (Fig. 8d).
For example, in Fig. 7d, cluster 0 was identified as acinar cells by the high expression of PRSS1. Clusters 2 and 7 were identified as alpha cells by the high expression of GCG. Cluster 3 was classified as ductal cells by the high expression of KRT19. Both COL1A1 and ESAM are highly expressed in cluster 4, suggesting mesenchymal or endothelial cells. Cluster 5 was classified as delta cells by the high expression of SST, while cluster 6 was classified as beta cells by the high expression of INS. The specific DEGs supporting each classification are detailed in Fig. 7a.
@@
To further validate the cell type assignments and clustering results, we cross-referenced the cluster-specific DEGs (see Fig.7a) with markers from the PanglaoDB cell type marker database (version 27 Mar 2020)[74]. The assignment of cell types was determined based on the overlap between the DEGs of each cluster and the cell type-specific markers in the PanglaoDB marker database, with a match score computed based on the proportion of matched markers (Table 7c). These results align well with our previous analysis based on marker gene distribution (Figure 7d). The only discrepancy lies in cluster 4. Matching the DEGs with PanglaoDB cell type markers suggests Pancreatic stellate cells (PSCs) or Granulosa cells, while the COL1A1 and ESAM expression plots indicate mesenchymal or endothelial cells. However, this apparent discrepancy can be explained by the fact that PSCs can differentiate into cells with endothelial-like properties [75], and granulosa cells can exhibit mesenchymal-like behaviors [76].
@@
In contrast, PCA clustering results (Fig.8) showed limited cluster separation. While some PCA clusters could be annotated consistently using both PanglaoDB markers and canonical markers, significant overlap remained among clusters. For example, PCA cluster 8, identified as beta cells by both the PanglaoDB markers and the canonical marker INS, is not well separated from cluster 5 (potentially delta cells, marked by SST expression). Similarly, cluster 7, identified as ductal cells by both the PanglaoDB markers and the canonical marker KRT19, is closely embedded with endothelial cells (ESAM). These results highlight CoMem-DIPHW’s improved ability to distinguish pancreatic cell types compared to PCA.
@@
Figure 7: Clustering Performance of CoMem-DIPHW on the Human Pancreas Dataset and Cell Type Annotation using Differentially Expressed Genes (DEGs) and Canonical Markers. (a) Top 10 DEGs per cluster identified by CoMem-DIPHW. For example, Cluster 0 (acinar cells) includes genes associated with digestive enzyme production, a primary function of pancreatic acinar cells, including CPA1, CPA2 [68], PLA2G1B [69], CTRB1 [70], and PRSS1 [71]. Clusters 2 and 7 (alpha cells) include markers GLS, CRYBA2, FAP, LOXL4, MAFB, and RGS4 (cluster 2) and LOXL4, PAX6, RGS4, and CRYBA2 (cluster 7). Cluster 3 (ductal cells) includes SLC4A4, ANXA4, KRT19, CFTR, CLDN1, and ONECUT2. Cluster 4, likely mesenchymal or endothelial cells, aligns with 9 of the top 10 DEGs found by CoMem-DIPHW, including COL4A2, COL4A1, SPARC, EMP1, IGFBP4, HTRA1, PXDN, LAMA4, and NID1. Cluster 5 (delta cells) includes GAD2, PCSK1, LEPR, and CASR, while Cluster 6 (beta cells) includes MAFA, ERO1LB, PDX1, IAPP, SYT13, HADH, and ADCYAP1. Cell type assignments are based on canonical marker expression patterns shown in (d). (b) Across-cluster average expression of cluster-specific DEGs. Violin plots show the distribution of average expression levels of these DEGs across all clusters. Strong clustering performance is indicated by high within-cluster expression of DEGs and low expression in other clusters. (c) Cell type annotation using the PanglaoDB marker database. Cell types are determined by the overlap between each cluster’s DEGs and PanglaoDB cell type-specific markers, with match scores computed based on the proportion of matched markers. (d) Expression of human pancreatic canonical markers in CoMem-DIPHW embeddings. UMAP visualizations show CoMem-DIPHW cell embeddings, with cluster IDs assigned by K-means clustering. Each subplot displays the expression of a canonical marker gene for a specific human pancreatic cell type. The cell-type-specific markers considered include: GCG (alpha cells), INS (beta cells), SST (delta cells), PPY (PP cells), PRSS1 (acinar cells), KRT19 (ductal cells), COL1A1 (mesenchymal cells), and ESAM (endothelial cells). Color intensity represents gene expression levels, with red hues indicating higher expression. Clusters show high expression of cell-type-specific canonical markers are assigned to the corresponding cell types. Overall, clustering quality is demonstrated through (b) high within-cluster DEG expression, (c) consistency between PanglaoDB-based annotation and clustering, and (d) distinct separation of cell types.
@@
Figure 8: Clustering Performance of PCA on the Human Pancreas Dataset and Cell Type Annotation using Differentially Expressed Genes (DEGs) and Canonical Markers. (a) Top 10 DEGs per cluster identified by PCA. (b) Across-cluster average expression of cluster-specific DEGs. Strong clustering performance is indicated by high expression of cluster-specific DEGs within their respective clusters and low expression in other clusters. (c) Cell type annotation using the PanglaoDB marker database. (d) Expression of human pancreatic canonical markers in PCA embeddings. The cell-type-specific markers considered include: GCG (alpha cells), INS (beta cells), SST (delta cells), PPY (PP cells), PRSS1 (acinar cells), KRT19 (ductal cells), COL1A1 (mesenchymal cells), and ESAM (endothelial cells). UMAP visualizations show PCA cell embeddings, with cluster IDs assigned by K-means clustering. The overall clustering of cell types is limited compared to CoMem-DIPHW clustering, as evidenced by the overlapping beta cells, delta cells and PP cells in (d) and the lack of distinct high expression patterns of the DEGs in (b).
@@
In addition to the human pancreas results presented here, we also analyzed mouse pancreas, human brain, and mouse brain datasets. For these datasets, we have included visualizations in the appendix that show the average expression of the top DEGs across clusters identified by different methods, which shows our proposed algorithm’s ability in effectively distinguishing cell clusters by their expression profiles. The supplements also contain tables listing the top DEGs for clusters identified by CoMem-DIPHW and PCA in each dataset for reference.